# Through the Lens of Split Vote: Exploring Disagreement, Difficulty and Calibration in Legal Case Outcome Classification

### v2

@@ Appendix G Debates in Legal Scholarship about Single Judge Dissenting Vote
Appendix H Judge’s Negligence Errors
Empirical legal scholarship identified time constraints leading to judge negligence errors in certain jurisdictions like the US Board of Veterans’ Appeals Ho et al. (2019) and Israeli Magistrate Courts Engel and Weinshall (2020). However, to the best our knowledge, there is no evidence has been documented in the ECtHR context to support judges’ careless behavior. Further, the Hungry Judge Effect, derived from a study, reveals that judges tend to issue harsher sentences just before lunch, presumably influenced by hunger Danziger et al. (2011). This implies that physiological states, like hunger, may affect rational decision-making. Nevertheless, a subsequent study Weinshall-Margel and Shapard (2011) suggests that the observed peak in favorable decisions after a meal break is likely an artifact of case presentation order, considering anticipated outcomes and duration. It does not serve as evidence that physiological states, such as hunger, impact judges’ rational decision-making.
Appendix I National Judge
In ECtHR, a national judge refers to the judge appointed from the respondent country of the case which was brought before the court. Like all other judges, national judges maintain independence and do not act as representatives of their respective governments. However, as previously noted in studies by legal schorlars, national judges have been observed to dissent more frequently in cases finding a violation of the Convention compared to their non-national counterparts Kuijer (1997); Helfer and Voeten (2020).Voeten 2008 provides evidence and explanations including for why the judges are considered policy-seekers and the answers is that judicial activism is driven by the political logic of European integration.
Appendix J Key Case
@@ Appendix I National Judge
Appendix K High Reputation Country
Previous legal and political science work indicates ECtHR Judges tend to have more split vote in cases where the defending country is regarded as High Reputation Country. Strong Democracies enjoy a high reputation in front of the judges while weak and new democracies only benefit from a low reputation implying that the ECHR should issue more demanding judgments against low reputation States than against high reputation States with the view not to damage its own reputation and ensure compliance with its decision Dothan (2014). We extracted the country information from the HUDOC metadata, namely Respondent State(s). Following in Chalkidis et al. 2022b, we group the countries according to the disproportion of violations between Eastern and Central European countries, and the rest of European Countries (Western European, Nordic, Mediterranean states).
Appendix L Vulnerability in ECtHR
The European Court of Human Rights (ECtHR) adapts convention standards to meet individual needs and thus to ensure effective human rights protection. Recognizing vulnerability is crucial for understanding unique needs and implementing targeted support systems to empower individuals in need. However, the concept of vulnerability remains undefined at the ECtHR, which can lead to divergent opinions on what qualifies as vulnerability from a legal point of view. Cases involving vulnerable applicants often deal with complex social-political issues related to the protection and rights of individuals who may face various challenges, such as victimization, migration, discrimination, reproductive health, unpopular views etc. The inclusion of vulnerability as a proxy variable is an example of how a human rights legal concept that is difficult to define can potentially give rise to diverging opinions on the bench. A comprehensive study of such concepts in ECtHR jurisprudence lies beyond the scope of this work. We direct the reader to Heri 2021 for systematic legal study and Xu et al. 2023a for comprehensive NLP research of Vulnerability in ECtHR.
Appendix M The impact of domain-specific pre-training on uncertainty representations
@@
Appendix N Ablation Experiment
@@
We also train hierachical LegalBERTs with the textified allegation information on the original dataset from Santosh et al. (2022) without corrected metadata. Results in Tab 7 show that model achieve better performance when fine-tuned on our dataset with corrected metadata.
Appendix O Rationale for Avoiding Implementation of Uncertainty Evaluation Methods
Within the deep learning community, uncertainty is often classified into two types: aleatoric and epistemic uncertainty. Epistemic (or model) uncertainty arises from a lack of knowledge about the best model, often exacerbated by out-of-distribution examples. Aleatoric uncertainty, on the other hand, stems from inherent ambiguity and can be considered as the variability in experiment outcomes Houlsby et al. (2011); Gal et al. (2016). As SV is due to inherent disagreement among judges, we regard it as aleatoic uncertainty.
Aleatoric uncertainty is typically quantified through metrics such as Entropy Gal et al. (2016) and Softmax Response (SR) Geifman and El-Yaniv (2017), as recognized in prior works Malinin and Gales (2018). However, studies show that these methods are based on predictive entropy, and actually measure total uncertainty, combining both epistemic and aleatoric uncertainty Van Amersfoort et al. (2020). Only when we have prior knowledge that either aleatoric or epistemic uncertainty is low , can we use predictive entropy as a suitable measure for the other type Mukhoti et al. (2023).
Given our small, label-imbalanced, and temporally shifted dataset, we choose to refrain from utilizing these methods to directly measure the aleatoric uncertainty of our cases. Instead, we opt to assess difficulty using Predictive Variance of Information (PVI), which measures predictive entropy and can be considered an equivalent for total uncertainty.
@@
Appendix P Details of the Models
P.1 Architecture of the Hierarchical Model
For the hierarchical variant of pre-trained BERT models, We use a greedy input packing strategy where we merge multiple paragraphs into one packet until it reaches the maximum of 512 tokens. We independently encode each packet of the input text using the pretrained model and obtain representations ($h_{[CLS]}$) for each packet. Then we apply a non-pretrained transformer encoder to make the packet representations context-aware. Fig 7 illustrates the detailed architecture of the hierarchical model.
P.2 Implementation Details
We use BERT "bert-base-uncased" Kenton and Toutanova (2019), and LegalBERT "nlpaueb/legal-bert-base-uncased" Chalkidis et al. (2020) from the Transformers Hub Wolf et al. (2020) as our backbone model.
Hyperparameter & Overfitting Measures: For the Hierarchical models, we employ a maximum sentence length of 128 and document length (number of sentences) of 80. The dropout rate in all layers is 0.1. We follow the hyperparameters from Chalkidis et al. 2022a with a batch size of 8 and learning_rate of 3e-3. We train models with the Adam optimizer up to 10 epochs. We use PyTorch Paszke et al. (2019) 2.0.1.
Appendix Q Soft Loss Function
@@
Appendix R Expected Calibration Error
Expected Calibration Error (ECE) is a measure of the difference between the predicted probabilities assigned by a model and the actual accuracy of those predictions. ECE is typically calculated by dividing the predicted probability space into a fixed number $M$ of intervals (or bins) and then computing the average absolute difference between the predicted and observed probabilities within each bin $B_{m}$ as follows:
@@
A lower ECE indicates better calibration, suggesting that the model’s predicted probabilities are more accurate reflections of the true likelihood. See Appendix for more about ECE.
Appendix S More Calibration Results
Tab 8 presents the calibration experiment results on both dev and test set. Soft-training improves the model’s COC performance (hm-F1) on SV cases for the development set, but lowers it for the test set. Two contributing factors account for this discrepancy: i) SV instances constitute less than 10% of the total dataset, potentially limiting the generalizability of improvements. ii) Given the temporal split of the dataset, distribution shifts in case-law may occur. For instance, SV cases could stabilize after a key legal decision. In other words, something that is legally controversial and uncertain in the training set time period may be settled law in the development and test dataset partitions. Our models may have overfit to the training set period. Fig 8 shows more figures comparing human uncertainty to model uncertainty.
@@
Table 8: COC performance (hm-F1), confidence-calibration (ECE), and human-calibration (DistCE) performance with std($\pm$) in dev and test set. "/": COC finetuned models on § 6.1; "soft": models fine-tuned with soft-loss; "TS": model after Temperature Scaling. We choose the temperature $t$ by searching a range of possible values for $t$ on the dev set. We noted that the chosen $t$ across three random seeds were consistently overly aggressive, with values of 5.5, 5.8, and 5.5. Calibration results with std; Results shown over 3 random seeds.

### v3

@@ Appendix F Debates in Legal Scholarship about Single Judge Dissenting Vote
Appendix G Judge’s Negligence Errors
Empirical legal scholarship identified time constraints leading to judge negligence errors in certain jurisdictions like the US Board of Veterans’ Appeals Ho et al. (2019) and Israeli Magistrate Courts Engel and Weinshall (2020). However, to the best our knowledge, there is no evidence has been documented in the ECtHR context for careless behavior by judges. The infamous Hungry Judge Effect, stemming from work in Danziger et al. (2011), suggested that judges tend to issue harsher sentences just before lunch, presumably influenced by hunger. However, a subsequent study Weinshall-Margel and Shapard (2011) argues that the observed peak in favorable decisions after a meal break is likely an artifact of case presentation order, considering anticipated outcomes and duration. Overall, this rebuts the notion that hunger impacts judges’ rational decision-making.
Appendix H National Judge
In the ECtHR context, a national judge refers to the judge appointed from the respondent country of the case which was brought before the court. Like all other judges, national judges maintain independence and do not act as representatives of their respective governments. However, as previously noted in studies by legal schorlars, national judges have been observed to dissent more frequently in cases finding a violation of the Convention compared to their non-national counterparts Kuijer (1997); Helfer and Voeten (2020).Voeten 2008 also provides evidence and explanations, including for why the judges are considered policy-seekers, and concludes that judicial activism is driven by the political logic of European integration.
Appendix I Key Case
@@
Appendix J High Reputation Country
Previous legal and political science work indicates ECtHR judges tend to have more split vote in cases where the defending country is regarded as a ‘High Reputation Country’. Strong democracies enjoy a high reputation in front of the judges while weak and new democracies only benefit from a low reputation implying that the ECtHR should issue more demanding judgments against low reputation states than against high reputation states, with the view not to damage its own reputation and ensure compliance with its decisions Dothan (2014). We extracted the country information from the HUDOC metadata, namely Respondent State(s). Following Chalkidis et al. 2022b, we group the countries according to the disproportion of violations between eastern and central European countries, and the rest of European countries (western European, Nordic, mediterranean states).
Appendix K Vulnerability in ECtHR
The ECtHR adapts convention standards to meet individual needs and to ensure effective human rights protection. Recognizing vulnerability is crucial for understanding unique needs and implementing targeted support systems. However, the concept of ‘vulnerability’ remains undefined by the court, which can lead to divergent opinions on what qualifies as vulnerability from a legal point of view. Cases involving vulnerable applicants often deal with complex social-political issues related to the protection and rights of individuals who may face various challenges, such as victimization, migration, discrimination, reproductive health, unpopular views etc. The inclusion of vulnerability as a proxy variable is an example of how a human rights legal concept that is difficult to define can potentially give rise to diverging opinions on the bench. A comprehensive study of such concepts in ECtHR jurisprudence lies beyond the scope of this work. We direct the reader to Heri 2021 for a systematic legal study and Xu et al. 2023a for comprehensive NLP research on vulnerability in the ECtHR context.
Appendix L The impact of domain-specific pre-training on uncertainty representations
@@
Appendix M Ablation Experiment
@@
We also train hierachical LegalBERTs with the textified allegation information on the original dataset from Santosh et al. (2022) without corrected metadata. Results in Tab 7 show that model achieves better performance when fine-tuned on our dataset with corrected metadata.
@@
Appendix N Details of the Models
N.1 Architecture of the Hierarchical Model
For the hierarchical variant of pre-trained BERT models, we use a greedy input packing strategy where we merge multiple paragraphs into one packet until it reaches the maximum of 512 tokens. We independently encode each packet of the input text using the pretrained model and obtain representations for each packet. Then we apply a non-pretrained transformer encoder to make the packet representations context-aware. Fig 7 illustrates the detailed architecture of the hierarchical model.
N.2 Implementation Details
We use BERT "bert-base-uncased" Kenton and Toutanova (2019), and LegalBERT "nlpaueb/legal-bert-base-uncased" Chalkidis et al. (2020) from the Transformers Hub Wolf et al. (2020) as our backbone models.
Hyperparameter & Overfitting Measures: For the hierarchical models, we employ a maximum sentence length of 128 and document length (number of sentences) of 80. The dropout rate in all layers is 0.1. We follow the hyperparameters from Chalkidis et al. 2022a with a batch size of 8 and learning_rate of 3e-3. We train models with the Adam optimizer for up to 10 epochs. We use PyTorch Paszke et al. (2019) 2.0.1.
Appendix O Soft Loss Function
@@
Appendix P Expected Calibration Error
Expected Calibration Error (ECE) is a measure of the difference between the predicted probabilities assigned by a model and the accuracy of those predictions. ECE is typically calculated by dividing the predicted probability space into a fixed number $M$ of intervals (or bins) and then computing the average absolute difference between the predicted and observed probabilities within each bin $B_{m}$ as follows:
@@
A lower ECE indicates better calibration, suggesting that the model’s predicted probabilities are more accurate reflections of the true likelihood.
Appendix Q More Calibration Results
Tab 8 presents the calibration experiment results on both dev and test set. Soft-training improves the model’s COC performance (hm-F1) on SV cases for the development set, but lowers it for the test set. Two contributing factors account for this discrepancy: i) SV instances constitute less than 10% of the total dataset, potentially limiting the generalizability of improvements. ii) Given the temporal split of the dataset, distribution shifts in case-law may occur. For instance, legal issues addressed in SV cases could stabilize after a key legal decision. In other words, something that is legally controversial and uncertain in the training set time period may be settled law in the development and test dataset partitions. Our models may have overfit to the training set period. Fig 8 shows more figures comparing human uncertainty to model uncertainty.
@@
Table 8: COC performance (hm-F1), confidence-calibration (ECE), and human-calibration (DistCE) performance with std($\pm$) in dev and test set. "/": COC finetuned models on § 6.1; "soft": models fine-tuned with soft-loss; "TS": model after Temperature Scaling. We choose the temperature $t$ by searching a range of possible values for $t$ on the dev set. We noted that the chosen $t$ across three random seeds were consistently overly aggressive, with values of 5.5, 5.8, and 5.5. Calibration results with standard deviation; Results shown over 3 random seeds.